OpenStudy (anonymous):

Uniqueness property? log2 2 (log base 2 can't write the little 2)

OpenStudy (anonymous):

\[\log_2(2)\]

OpenStudy (anonymous):

yes sorry I don't know how to make the small 2

OpenStudy (anonymous):

solve \[2^x=2\]for \(x\) in your head

OpenStudy (anonymous):

\[2^1=2\]

OpenStudy (anonymous):

of course that makes \[\log_2(2)=1\]

OpenStudy (anonymous):

it is always true that \[\log_b(b^x)=x\]

OpenStudy (anonymous):

so my answer is 1?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

and if you wanted \[\log_2(8)\] you would solve \[2^x=8\] since \[2^3=8\] that means \[\log_2(8)=3\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[ \log_{4} 16 \] for this would it be \[4^2\] = 16

OpenStudy (anonymous):

yes, and therefore \[\log_4(16)=2\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw